Problem: From the information below,write an equation and then solve the equation.
My mother is three times as old as my sister. If the difference between their ages is 22 years, find both of their ages.

Answer by Earlsdon(6294)

Let M = mother's age and S = sister's age.
From the problem description, you can write:
M = 3S Mother's age is three times sister's age.
M-S = 22 The difference in their ages is 22. In this equation, substitute M = 3S.
3S-S = 22
2S = 22
S = 11
M = 3(11)
M = 33
Mother's age is 33 years.
Sister's age is 11 years.